# Post-hoc Reward Calibration: A Case Study on Length Bias

### v1

@@ Abstract
Reinforcement Learning from Human Feedback aligns the outputs of Large Language Models with human values and preferences. Central to this process is the reward model (RM), which translates human feedback into training signals for optimising LLM behaviour. However, RMs can develop biases by exploiting spurious correlations in their training data, such as favouring outputs based on length or style rather than true quality. These biases can lead to incorrect output rankings, sub-optimal model evaluations, and the amplification of undesirable behaviours in LLMs alignment. This paper addresses the challenge of correcting such biases without additional data and training, introducing the concept of Post-hoc Reward Calibration. We first propose an intuitive approach to estimate the bias term and, thus, remove it to approximate the underlying true reward. We then extend the approach to a more general and robust form with the Locally Weighted Regression. Focusing on the prevalent length bias, we validate our proposed approaches across three experimental settings, demonstrating consistent improvements: (1) a 3.11 average performance gain across 33 reward models on the RewardBench dataset; (2) enhanced alignment of RM rankings with GPT-4 evaluations and human preferences based on the AlpacaEval benchmark; and (3) improved Length-Controlled win rate of the RLHF process in multiple LLM–RM combinations. Our method is computationally efficient and generalisable to other types of bias and RMs, offering a scalable and robust solution for mitigating biases in LLM alignment. Our code and results are available at https://github.com/ZeroYuHuang/Reward-Calibration.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Kaufmann et al., 2023) is driving the success of Large Language Models (LLMs) (OpenAI, 2023; Anil et al., 2023; Dubey et al., 2024). By integrating human preferences into the training loop, RLHF aligns models with desired behaviours and human values (Chakraborty et al., 2024), enabling LLMs to generate content that is conceptually appropriate (Sun et al., 2023), ethically sound (Hendrycks et al., 2021a; Dai et al., 2024), and reflecting nuanced human judgements (Wu et al., 2023). A key component in this process is the reward model (RM), which translates qualitative human feedback into a signal for optimising LLMs. The primary purpose of RM training is to approximate human preferences, allowing it to evaluate the content generated by LLMs by assigning scores. These scores (rewards) are then used to optimize the model, shaping its behaviour to better meet human expectations.
@@
However, there is a risk that an RM may exploit spurious correlations in the training data that do not genuinely reflect the output quality (Stiennon et al., 2020; Gao et al., 2023; Singhal et al., 2023; Pang et al., 2023; Lambert & Calandra, 2023). For instance, an RM might favour outputs based on length (Park et al., 2024a) or style (Lambert et al., 2024; Dubois et al., 2023a), leading to biased assessments that do not accurately represent human preferences. The involvement of a biased RM in developing LLMs can have significant drawbacks. Firstly, it may lead to incorrect rankings of outputs, favouring content that aligns with learned biases over genuinely high-quality outputs. This issue is particularly problematic when RMs are used to rank different models during LLM evaluations, as a biased RM could produce misleading results, endorsing suboptimal models and impeding progress in model development. Moreover, when biased RMs are employed in RLHF, these biases can be further amplified (Zheng et al., 2023b), with the model learning to exploit these biases rather than improving genuine task performance—a phenomenon known as reward hacking.
@@
Recent advancements introduce various strategies to mitigate these biases and alleviate reward hacking during different stages of RLHF, from data handling (Park et al., 2024a) and reward model engineering (Coste et al., 2024) to the final reinforcement learning phase (Chen et al., 2024). Most of these approaches, however, either require additional data collection, re-engineering of the RM, or modifications to reinforcement learning algorithms, which raises an intriguing question: is it possible to correct or mitigate biases in reward signals in a training-free manner? Specifically, given a set of scored outputs from the RM and a specific bias that the model may exhibit, how can we calibrate these reward signals to reflect the true quality of the outputs more accurately?
@@
This paper makes an initial attempt at answering this question by framing it as Post-hoc Reward Calibration. Specifically, for a particular characteristic of the output (e.g., the length), we assume the biased reward from the RM can be decomposed into the sum of (1) the underlying calibrated reward and (2) a bias term that solely depends on that characteristic. Under certain assumptions, we propose to use the local average reward to provide an estimation for the bias term, which can then be removed, thereby approximating the true reward. We then extend it to be more general and robust using Locally Weighted Regression for bias estimation. Among different biases (Zhang et al., 2024b), a particularly prevalent one is the sequence length. Prior works shown that models generate significantly longer responses after RLHF without improving the genuine quality of the content (Kabir et al., 2023; Wu et al., 2023; 2024). Singhal et al. (2023) point out that even a purely length-based reward can reproduce most downstream RLHF improvements of existing RMs, highlighting that current RMs fail to convincingly outperform simple heuristics and may be strongly biased by the length. The issue is also evident in powerful proprietary models like GPT-4, which is increasingly used as a judge, resulting in LLM evaluations that favour more verbose responses. Focusing on length bias, we validate our method in three settings and observe consistent improvements: (1) Benchmark performance: a 3.11 performance gain averaged over 33 RMs on the RewardBench; (2) RM as LLMs evaluators: based on AlpacaEval, we utilise 8 open-source RMs to rank 184 LLMs. After applying our calibration, the rankings correlate better with GPT-4 evaluations and human preferences. (3) RM for LLMs alignment: we assess calibrated rewards for RLHF process. Testing with four LLM–RM configurations, we observe consistent improvements in AlpacaEval2 Length-controlled win rates.
@@
In addition to being computationally efficient (e.g., calibrating over 300k samples takes only 30 seconds with a single CPU), we empirically reveal that our method can generalise to other biases, such as the presence of markdown-style features in the outputs, and to pairwise GPT4-as-Judge models. Importantly, the calibration effect is strongly correlated with the RM’s preference for specific properties: for strongly biased RMs, the calibration yields greater improvements, while for weakly biased RMs, it only slightly alters its rewarding behaviour. These findings suggest that our proposed Post-hoc Reward Calibration method enhances the reliability of RMs and offers a practical, scalable solution for mitigating biases in RLHF. This approach could be instrumental in advancing RLHF methodologies and developing more robust and aligned LLMs.
@@ 2 Related works
In practice, RMs are often implemented by adding an extra linear layer to LLMs, which projects the representations of the last layers of LLMs to a scalar as the reward. There are also two other types of RMs: (1) DPO-based: Models optimised with Direct Preference Optimisation (DPO) (Rafailov et al., 2023) could function as an implicit RM. (2) LLM-as-Judge: A generative LLM can also be prompted to function as an RM to generate feedback, usually consisting of a piece of chain-of-thought reasoning and a final score representing the overall quality (Zheng et al., 2023a; Kwon et al., 2023; Kim et al., 2024). In this paper, we validate the effectiveness of our methods across all three types of judge models. Specifically, we employ our method to calibrate BT-based and DPO-based RMs on the RewardBench dataset, and to calibrate the pairwise judge of GPT-4 with the AlpacaEval benchmark.
@@
Recent works explore the bias mitigation, also termed as mitigate reward over-optimisation, for the RLHF process mainly from the perspective of (1) Data handling:  Park et al. (2024a) identifies six types of inherent bias in various judge models and employ LLMs to generate the de-biasing dataset for further fine-tuning.  Zhu et al. (2024) propose to iteratively replace hard labels in data with soft labels during RM training, while  Rita et al. (2024) rely on human expert demonstration to recalibrate the reward objective optimized with Reinforcement Learning. (2) Reward Model Engineering:  Coste et al. (2024); Eisenstein et al. (2023); Sun et al. (2024); Zhang et al. (2024a) implement the reward ensemble;  Ramé et al. (2024) first fine-tune multiple RMs and then average them in the weight space;  Dai et al. (2024) decouple the reward model training for helpfulness and harmlessness and  Chen et al. (2024) disentangle the length signal and the quality signal into two different reward heads.  Shen et al. (2023); Wang et al. (2024); Quan (2024) exploit the idea of Mixture-of-Experts (Qiu et al., 2024a; b) in RM training to capture different facets of human preference to avoid the potential conflict between them; (3) Reinforcement Learning methods: Another research line  (Moskovitz et al., 2024; Sun et al., 2024; Zhou et al., 2024; Zhai et al., 2024) modifies the algorithm or optimisation objective during reinforcement learning with  Park et al. (2024b); Zhu et al. (2024); Lu et al. (2024) specifically focusing on the length bias. Our work takes a different approach from previous methods by employing post-hoc reward calibration, i.e., using only a batch of scored prompt-response examples to mitigate the reward model’s bias, without intervening the preference data collection, RM training, and the RLHF phase.
@@ 3.1 Problem Statement: Reward Calibration
${\rm c}:$ ${\mathbb{X}}$ $\mapsto \displaystyle\quad{\mathbb{R}}$
@@
We make the following hypothesis about the decomposition in Eq. 3. (1) Independence of the biased characteristic: consider a general reward modeling dataset composed of prompts and corresponding responses pairs $\mathcal{D}=\{(q^{i},y^{i}_{1},y^{i}_{2})\}_{i=1}^{N}=\{(x_{1}^{i},x_{2}^{i})\}_{i=1}^{N}$, the expectation of the underlying gold reward margin should be 0 and is independent of the biased characteristic $c$,
@@
(2) Sufficient Density: the function values of the characteristic function $c$ is assumed to be sufficiently dense in its range $c({\mathbb{X}})$. (3) Lipschitz Continuity: the bias term $b_{c}^{\theta}$ is a slow-varying function of characteristic $c$, which means if pairs $x_{1}$ and $x_{2}$ are close to each other regarding the characteristic $c$, their corresponding bias term in Eq. 3 should be close as well. Mathematically, given $c(x_{0})$ and a neighbourhood $\mathbb{V}(x_{0})=\{c(x)\mid\lvert c(x)-c(x_{0})\rvert<\epsilon\}$ of $x_{0}$, for all $c(x_{1}),c(x_{2})\in\mathbb{V}$, there exists a constant $K$ such that $\lvert b_{c}^{\theta}(c(x_{1}))-b_{c}^{\theta}(c(x_{1}))\rvert<K\lvert c(x)-c(x_{0})\rvert$;
@@ 3.2 Bias Estimation
Under the assumption of Lipschitz continuity and sufficient density, we can obtain a simple calibrated reward margin described in the Eq. 6:
@@
where $\hat{r}_{\theta}(c(x_{1}))$ and $\hat{r}_{\theta}(c(x_{2}))$ are the LWR prediction for the characteristic $c(x_{1})$ and $c(x_{2})$, respectively, given the dataset $D_{c}$. To determine the preference, we use the sign of reward margin between $x_{1}$ and $x_{2}$: $\operatorname{sgn}(\hat{\Delta}^{*}_{r_{\theta}}(x_{1},x_{2}))$. In practice, we employ the Robust Locally Estimated Scatterplot Smoothing (LOWESS), a robust version of LWR. Based one LWR, it defines different weights $\delta_{j}$ for the point $(c(x_{j}),r_{\theta}(x_{j}))$ according to the size of residual $r_{\theta}(x_{j})-\hat{r}_{\theta}(c(x_{j}))$. It assigns smaller weights to data points with large residuals and large weights to those with small residuals. New fitted values are computed using LWR again but with $w_{j}$ replaced by $\delta_{j}w_{j}$. The computation of new weights and fitted values is repeated several times to converge. The full algorithm of using LOWESS for reward calibration is described with Algorithm 1. We directly employ off-the-shelf library statsmodels.api and pass the dataset $D_{c}$ to it for our implementation.
@@ 4 Experiments
We test the performance of the proposed reward calibration method in three settings: (1) RM Benchmark performance: we test the calibrated RM’s performance on the RewardBench dataset (Lambert et al., 2024). (2) RM As LLMs Evaluator: Based on the AlpacaEval benchmark (Dubois et al., 2023a), we employ RMs to label the output of different LLMs from AlpacaEval leaderboard and produce a ranking of LLMs. Then, we measure the correlation between the RM’s ranking and GPT-4 ranking (AlpacaEval2) or human preference ranking (ChatbotArena). (3) RM for LLMs alignment: We utilise the calibrated rewards rather than the original reward to label the preference data, and then we conduct the Direct Preference Optimization (Rafailov et al., 2023) with the labelled preference pair. We evaluate the aligned model’s performance using AlpacaEval2 and eight popular benchmarks. In all three settings, we are given a group of data points $D_{|x|}=\{x_{i},r_{\theta}(x_{i})\}$ for calibration. Because LOWESS requires calculating the residual for every fitted data point, we employ the entire $D_{|x|}$ for regression. Therefore, for RewardBench, there are 2,985 test data points, so we employ all 5,970 data points to perform the regression (every test data is a pair of responses for one prompt). Regarding AlpacaEval, there are 184 LLMs and 805 instructions for each model, resulting in 151k samples used for calibration. For LLMs alignment, five responses are generated from the model for about 60k instructions, leading to about 300k samples used for bias estimation. The proposed method is computation-efficient, and calibrating 300k samples only takes 30 seconds using a single CPU.
@@
For all three experimental settings, we try to compare the following algorithms: (1) Original reward; (2) Length penalty: a broadly adopted approach to mitigate the length bias in reward models. It adds a penalty term to the original reward (Singhal et al., 2023; Park et al., 2024b; Liu et al., 2024; Dong et al., 2024), formalised as $\hat{r^{*}}(x)=r_{\theta}(x)-\alpha\times|x|$. Following Dong et al. (2024), we use the character length of each output and set $\alpha=0.001$. (3) RC-Mean: our proposed uniform averaging calibration method described in Eq. 6. (4) RC-LWR: our proposed Locally Weighted Regression calibration method shown in Eq. 7. (5) RC-LWR-Penalty: to see whether our method could be applied on top of the penalty-based methods, we first apply the length penalty to the original reward and then use the RC-LWR. All hyper-parameters related to RC-Mean, RC-LWR and RC-LWR-Penalty are presented in Appendix B.

### v2

@@ Abstract
Reinforcement Learning from Human Feedback aligns the outputs of Large Language Models with human values and preferences. Central to this process is the reward model (RM), which translates human feedback into training signals for optimising LLM behaviour. However, RMs can develop biases by exploiting spurious correlations in their training data, such as favouring outputs based on length or style rather than true quality. These biases can lead to incorrect output rankings, sub-optimal model evaluations, and the amplification of undesirable behaviours in LLMs’ alignment. This paper addresses the challenge of correcting such biases without additional data and training, introducing the concept of Post-hoc Reward Calibration. We first propose to use the local average reward to estimate the bias term and, thus, remove it to approximate the underlying true reward. We then extend the approach to a more general and robust form with the Locally Weighted Regression. Focusing on the prevalent length bias, we validate our proposed approaches across three experimental settings, demonstrating consistent improvements: (1) a 3.11 average performance gain across 33 reward models on the RewardBench dataset; (2) improved agreement of RM produced rankings with GPT-4 evaluations and human preferences based on the AlpacaEval benchmark; and (3) improved Length-Controlled win rate (Dubois et al., 2024) of the RLHF process in multiple LLM–RM combinations. According to our experiments, our method is computationally efficient and generalisable to other types of bias and RMs, offering a scalable and robust solution for mitigating biases in LLM alignment and evaluation.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022; Kaufmann et al., 2023) is driving the success of Large Language Models (LLMs) (OpenAI, 2023; Anil et al., 2023; Dubey et al., 2024). By integrating human preferences into the training loop, RLHF aligns models with desired behaviours and human values (Chakraborty et al., 2024), enabling LLMs to generate content that is conceptually appropriate (Sun et al., 2023), ethically sound (Hendrycks et al., 2021a; Dai et al., 2024), and reflecting nuanced human judgements (Wu et al., 2023). A key component in this process is the reward model (RM), which translates qualitative human feedback into a signal for optimising LLMs. The primary purpose of RM training is to approximate human preferences, allowing it to evaluate the content generated by LLMs by assigning scores. These scores (rewards) are then used to optimize the LLM, shaping its behaviour to better meet human expectations.
@@
However, there is a risk that an RM may exploit spurious correlations in the training data that do not genuinely reflect the output quality (Stiennon et al., 2020; Gao et al., 2023; Singhal et al., 2023; Pang et al., 2023; Lambert & Calandra, 2023). For instance, an RM might favour outputs based on length (Park et al., 2024a) or style (Lambert et al., 2024; Dubois et al., 2023a), leading to biased assessments that do not accurately represent human preferences. The involvement of a biased RM in developing LLMs can have significant drawbacks. Firstly, it may lead to incorrect rankings of outputs, favouring content that aligns with learned biases over genuinely high-quality outputs. This issue is particularly problematic when RMs are used to rank different LLMs during evaluations, as a biased RM could produce misleading results, endorsing suboptimal models and impeding progress in model development. Moreover, when biased RMs are employed in RLHF, these biases can be further amplified (Zheng et al., 2023b), with the LLM learning to exploit these biases rather than improving genuine task performance—a phenomenon known as reward hacking.
@@
Recent advancements introduce various strategies to mitigate these biases and alleviate reward hacking during different stages of RLHF, from data handling (Park et al., 2024a) and reward model engineering (Coste et al., 2024) to the final reinforcement learning phase (Chen et al., 2024b). Most of these approaches, however, either require additional data collection, re-engineering of the RM, or modifications to reinforcement learning algorithms, which raises an intriguing question: is it possible to correct or mitigate biases in reward signals in a training-free manner? Specifically, given a set of scored outputs from the RM and a specific bias that the RM may exhibit, how can we calibrate these reward signals to reflect the true quality of the outputs more accurately?
@@
This paper makes an initial attempt at answering this question by framing it as Post-hoc Reward Calibration. Specifically, for a particular characteristic of the output (e.g., the length), we assume the biased reward from the RM can be decomposed into the sum of (1) the underlying calibrated reward and (2) a bias term that solely depends on that characteristic. Under certain assumptions, we propose to use the local average reward to provide an estimation for the bias term, which can then be removed, thereby approximating the true reward. We then extend it to be more general and robust using Locally Weighted Regression for bias estimation. Among different biases (Zhang et al., 2024b), a particularly prevalent one is the sequence length. Prior works showed that LLMs generate significantly longer responses after RLHF without improving the genuine quality of the content (Kabir et al., 2023; Wu et al., 2023; 2024b). Singhal et al. (2023) point out that even a purely length-based reward can reproduce most downstream RLHF improvements of existing RMs, highlighting that current RMs fail to convincingly outperform simple heuristics and may be strongly biased by length. The issue is also evident in powerful proprietary models like GPT-4, which is increasingly used as a judge, resulting in LLM evaluations that favour more verbose responses. Focusing on length bias, we validate our method in three settings and observe consistent improvements: (1) Benchmark performance: a 3.11 performance gain averaged over 33 RMs on the RewardBench; (2) RM as LLMs evaluators: based on AlpacaEval, we utilise 8 open-source RMs to rank 184 LLMs. After applying our calibration, the rankings correlate better with GPT-4 evaluations and human preferences. (3) RM for LLMs’ alignment: we assess calibrated rewards for RLHF process. Testing with four LLM–RM configurations, we observe consistent improvements in AlpacaEval2 Length-controlled win rates.
@@
In addition to being computationally efficient (e.g., calibrating over 300k samples takes only 30 seconds with a single CPU), we empirically reveal that our method can generalise to other quantifiable biases, such as the presence of markdown-style features in the outputs, and to pairwise GPT4-as-Judge models. Importantly, the calibration effect is strongly correlated with the RM’s preference for specific properties: for strongly biased RMs, the calibration yields greater improvements, while for weakly biased RMs, it only slightly alters its rewarding behaviour. These findings suggest that our proposed Post-hoc Reward Calibration method enhances the reliability of RMs and offers a practical, scalable solution for mitigating biases in RLHF. This approach could be instrumental in advancing RLHF methodologies and developing more robust and aligned LLMs.
@@ 2 Related works
In practice, RMs are often implemented by adding an extra linear layer to LLMs, which projects the representations of the last layers of LLMs to a scalar as the reward. There are also two other types of RMs: (1) DPO-based: LLMs optimised with Direct Preference Optimisation (DPO) (Rafailov et al., 2023) could function as an implicit RM. (2) LLM-as-Judge: A generative LLM can also be prompted to function as an RM to generate feedback, usually consisting of a piece of chain-of-thought reasoning and a final score representing the overall quality (Zheng et al., 2023a; Kwon et al., 2023; Kim et al., 2024). In this paper, we validate the effectiveness of our methods across all three types of judge models. Specifically, we employ our method to calibrate BT-based and DPO-based RMs on the RewardBench dataset, and to calibrate the pairwise judge of GPT-4 with the AlpacaEval benchmark.
@@
Recent works explore the bias mitigation, also termed as mitigate reward over-optimisation, for the RLHF process mainly from the perspective of (1) Data handling:  Park et al. (2024a) identifies six types of inherent bias in various judge models and employ LLMs to generate the de-biasing dataset for further fine-tuning.  Zhu et al. (2024) propose to iteratively replace hard labels in data with soft labels during RM training, while  Rita et al. (2024) rely on human expert demonstration to recalibrate the reward objective optimised with Reinforcement Learning. (2) Reward Model Engineering:  Coste et al. (2024); Eisenstein et al. (2023); Sun et al. (2024); Zhang et al. (2024a) implement the reward ensemble;  Ramé et al. (2024) first fine-tune multiple RMs and then average them in the weight space;  Dai et al. (2024) decouple the RM training for helpfulness and harmlessness and  Chen et al. (2024b) disentangle the length signal and the quality signal into two different reward heads.  Shen et al. (2023); Wang et al. (2024); Quan (2024) exploit the idea of Mixture-of-Experts (Qiu et al., 2024a; b) in RM training to capture different facets of human preference to avoid the potential conflict between them; (3) Reinforcement Learning methods: Another research line  (Moskovitz et al., 2024; Sun et al., 2024; Zhou et al., 2024; Zhai et al., 2024) modifies the algorithm or optimisation objective during reinforcement learning with  Park et al. (2024b); Zhu et al. (2024); Lu et al. (2024) specifically focusing on the length, one of the most consequential sources of bias (More discussions on length bias are in Appendix. A). Our work takes a different approach from previous methods by employing post-hoc reward calibration, i.e., using only a batch of scored prompt-response examples to mitigate the RM’s bias, without intervening the preference data collection, RM training, and the RLHF phase.
@@ 3.1 Problem Statement: Reward Calibration
${\rm c}:$ ${\mathbb{X}}$ $\mapsto$ $\quad{\mathbb{R}}$
@@
We make the following hypothesis about the decomposition in Eq. 3. (1) Independence of the biased characteristic: consider a general reward modelling dataset composed of prompts and corresponding responses pairs $\mathcal{D}=\{(q^{i},y^{i}_{1},y^{i}_{2})\}_{i=1}^{N}=\{(x_{1}^{i},x_{2}^{i})\}_{i=1}^{N}$, the expectation of the underlying gold reward margin should be 0 and is independent of the biased characteristic $c$,
@@
(2) Sufficient Density: the function values of the characteristic function $c$ is assumed to be sufficiently dense in its range $c({\mathbb{X}})$. (3) Lipschitz Continuity: the bias term $b_{c}^{\theta}$ is a slow-varying function of characteristic $c$, which means if pairs $x_{1}$ and $x_{2}$ are close to each other regarding the characteristic $c$, their corresponding bias term in Eq. 3 should be close as well. Mathematically, given $c(x_{0})$ and a neighbourhood $\mathbb{V}(x_{0})=\{c(x)\mid\lvert c(x)-c(x_{0})\rvert<\epsilon\}$ of $x_{0}$, for all $c(x_{1}),c(x_{2})\in\mathbb{V}$, there exists a constant $K$ such that $\lvert b_{c}^{\theta}(c(x_{1}))-b_{c}^{\theta}(c(x_{1}))\rvert<K\lvert c(x)-c(x_{0})\rvert$.
@@ 3.2 Bias Estimation
Under the assumption of Lipschitz continuity and sufficient density, we can obtain a simple calibrated reward margin described in Eq. 6:
@@
where $\hat{r}_{\theta}(c(x_{1}))$ and $\hat{r}_{\theta}(c(x_{2}))$ are the LWR prediction for the characteristic $c(x_{1})$ and $c(x_{2})$, respectively, given the dataset $D_{c}$. To determine the preference, we use the sign of reward margin between $x_{1}$ and $x_{2}$: $\operatorname{sgn}(\hat{\Delta}^{*}_{r_{\theta}}(x_{1},x_{2}))$. In practice, we employ the Robust Locally Estimated Scatterplot Smoothing (LOWESS), a robust version of LWR. Based on LWR, it defines different weights $\delta_{j}$ for the point $(c(x_{j}),r_{\theta}(x_{j}))$ according to the size of residual $r_{\theta}(x_{j})-\hat{r}_{\theta}(c(x_{j}))$. It assigns smaller weights to data points with large residuals and large weights to those with small residuals. New fitted values are computed using LWR again but with $w_{j}$ replaced by $\delta_{j}w_{j}$. The computation of new weights and fitted values is repeated several times to converge. The full algorithm of using LOWESS for reward calibration is described with Appendix B Algorithm 1. We directly employ off-the-shelf library statsmodels.api and pass the dataset $D_{c}$ to it for our implementation.
@@ 4 Experiments
One key advantage of our proposed Post-hoc reward calibration is that it does not need any extra data or training and thus can be seamlessly applied to various scenarios. So, we test the proposed method in three settings: (1) RM Benchmark performance: we test the calibrated RM’s performance on the RewardBench dataset (Lambert et al., 2024). (2) RM As LLMs Evaluator: Another critical application for RM is to serve as a judge to replace human evaluation when evaluating LLMs. We thus also explore whether the calibrated RM could provide a better evaluation. Based on the AlpacaEval benchmark (Dubois et al., 2023a), we employ RMs to label the output of different LLMs from AlpacaEval leaderboard and produce a ranking of LLMs. Then, we measure the correlation between the RM’s ranking and GPT-4 ranking or human preference ranking. (3) RM for LLMs’ alignment: Finally, as discussed in our introduction, employing a biased RM risks amplifying the potential bias and thus leads to unexpected behaviours. So, we explore how our proposed calibration method affects the LLM alignment results. Specifically, we use the calibrated rewards to label the preference data, and then conduct Direct Preference Optimization (Rafailov et al., 2023) with the labelled preference pair. We evaluate the aligned LLM’s performance using AlpacaEval2 and eight popular benchmarks. In all three settings, we are given a group of data points $D_{|x|}=\{x_{i},r_{\theta}(x_{i})\}$ for calibration. Because LOWESS requires calculating the residual for every fitted data point, we employ the entire $D_{|x|}$ for regression. Therefore, for RewardBench, there are 2,985 test data points, so we employ all 5,970 data points to perform the regression (every test data is a pair of responses for one prompt). Regarding AlpacaEval, there are 184 LLMs and 805 instructions for each LLM, resulting in 151k samples used for calibration. For LLMs’ alignment, five responses are generated from the LLM for about 60k instructions, leading to about 300k samples used for bias estimation. The proposed method is computation-efficient, and calibrating 300k samples only takes 30 seconds using a single CPU.
@@
For all three experimental settings, we try to compare the following algorithms: (1) Original reward; (2) Length penalty: a broadly adopted approach to mitigate the length bias in reward models. It adds a penalty term to the original reward (Singhal et al., 2023; Park et al., 2024b; Liu et al., 2024a; Dong et al., 2024), formalised as $\hat{r^{*}}(x)=r_{\theta}(x)-\alpha\times|x|$. Following Dong et al. (2024), we use the character length of each output and set $\alpha=0.001$. (3) RC-Mean: our proposed uniform averaging calibration method described in Eq. 6. (4) RC-LWR: our proposed Locally Weighted Regression calibration method shown in Eq. 7. (5) RC-LWR-Penalty: to see whether our method could be applied on top of the penalty-based methods, we first apply the length penalty to the original reward and then use the RC-LWR. All hyper-parameters related to RC-Mean, RC-LWR and RC-LWR-Penalty are presented in Appendix C.